OpenStudy (liizzyliizz):

How do solve for the cubed root of 32. .-. For some reason I can't do it.

OpenStudy (venomblast):

first you need to find the a perfect cube root and reduce the radicals. so it would be \[\sqrt[3]{8}\] and \[\sqrt[3]{4}\]. you can simply the \[\sqrt[3]{8}\] to 2. so the answer would be \[2\sqrt[3]{4}\]